OpenStudy (anonymous):

Find an equation of the line. Perpendicular to y=3x+5; through (1,0). Write the equation in Standard Form.

OpenStudy (anonymous):

Your original equation is given is y = mx + b form. In this form m = 3 is your slope. Perpendicular lines have opposite reciprocal slopes (not a complicated as it sounds). Basically flip the slope over and multiply by -1. So the new slope becomes -1/3. Now you have a slope and a point, begging for point slope form: y - y1 = m(x - x1). From there, you need to convert to standard form.

OpenStudy (anonymous):

so does y=-1/3 ?

OpenStudy (anonymous):

No, your slope is -1/3.

OpenStudy (anonymous):

\[y-0=-\frac{1}{3}(x-1)\] is the point slope form or just \[y=-\frac{1}{3}(x-1)\] multiply out if you like

OpenStudy (anonymous):

ooh standard form, i should learn to read \[y=-\frac{1}{3}(x-1)\] multiply both sides by \(-3\) and get \[-3y=x-1\]

OpenStudy (anonymous):

which you can then rewrite as \[x+3y=1\] if you like

OpenStudy (anonymous):

i'm very confused ;_;